# What-if Analysis for Business Users: Current Practices and Future Opportunities

### v3

@@ 7.2.3. Streamlines Decision-Making and Enhances Teamwork
Almost all participants agreed that their existing analysis processes could not be replicated easily to justify their decision actions. For instance, multiple participants commented that their current decision actions were based on “trial and error” (P3, P14, P19, P21), which they often deemed to be inefficient and unpredictable, making it challenging to rely on. Recognizing the limitations of their current practices, participants saw the potential of WIA to streamline decision processes, making them more accessible, reproducible, and actionable. This consequently allowed for “consistent” (P3, P10) (Kandel et al., 2012) and “smoother” (P22) decision-making. Further, in addition to WIA techniques helping them in their own analysis, participants also added that it will help facilitate collaborative analyses for decision-making envisioning a scenario where stakeholders could engage in decision reasoning collectively. For example, one participant P4 mentioned how they could “take [this] to their job today without the support of IT… also give it to other team members”. Therefore, participants explained that advanced analytics techniques would help them make robust decisions as business users could get external perspectives and benefits from others’ experiences and expertise through standardized and collaborative analysis processes.
@@ 7.3. Opportunities in Enhancing the Use of WIA for Making Data-Driven Decisions
7.3.1. Currently Inaccessible to Business Users
Many participants like P1, P7, P8 mentioned that even though they had heard of features similar to some representative WIA techniques in existing commercial BI tools (e.g., key influencers in Power BI, importance and correlation analysis in Einstein, etc.), they had “no idea or not know too much [about them] or barely used them”. Upon further discussion, participants revealed that high costs associated with such tools rendered them inaccessible. Additionally participants voiced concerns about the complexity of these tools, which typically offered a plethora of models and parameters to navigate, or deemed them to be too general and inefficient to use for their own use cases. The overwhelming number of features too made it challenging for users to effectively utilize the tools, exacerbated by automatically generated results with a lack of clear explanations for them to draw conclusions from. Moreover, participants cautioned the requirement of technical expertise to utilize these tools effectively, a skill set that many business users did not possess. Furthermore, business users also lacked the time and incentives to learn or to keep up with such technologies, which are rapidly evolving.
Despite their enthusiasm for adopting WIA in their work, it seems that tools are not currently designed to meet their needs. The vast space of possible scenarios seems to be a major problem in particular. P15, for instance, frustratingly mentioned how they have “…to [currently] commit to just solving the first thing… that is going to be the most bang for the buck…. [because]… have 900 problems happening all at once…”. Such expressions inform us that business users have no time to experiment with all possible scenarios but need flexibility to quickly identify the first set of actionable decisions that yield favorable business outcomes and play with them to account for constraints, a concept akin to “satisficing” (Simon, 1960), suggesting that more complex analyses atop WIA, like recommendations or ranking of WIA scenarios, could help.
@@ 7.3.2. Difficulties in Interpreting and Understanding Predictions
While participants expressed enthusiasm for utilizing WIA techniques for decision-making, they were clear that such analytics alone cannot make the decisions. Their domain expertise remains critical in strategizing decisions (Wang et al., 2019). For instance, P4 pointed to how “common human understanding may not translate to the model without [their] external involvement”, giving an example where an increase in unit price beyond a threshold will decrease demand, thereby decreasing sales rather than increasing it, a nuance that may not be fully captured by models. Participants emphasized the need for transparency in how predictions were generated and wanted clarification on when and to what extent to adopt predictions. For instance, they wanted to know how to handle “unrealistic [predictions]” (P5, P6, P22), or when “so many factors [or decision paths]” (P2) are shown to them, and how to “[include] factors that cannot be captured in the data” (P2, P18).
Another recurring challenge was interpreting the accuracy of predictions (low v/s fair v/s high) and translating this into confidence in next steps (Elhamdadi et al., 2022). For example, should low-scored models be treated as worst-case scenarios and perform additional manual analysis to make decisions? Or should they take medium scored models as baseline and expect risks in lower scored models? While senior, more experienced business users understood that predictions should be taken with caution, they emphasized the importance of training junior users not to follow predictions blindly. Addressing these concerns requires WIA tools to foster deeper transparency and provide interpretable predictions that can help enhance business users’ trust in the predictions.
@@ 7.3.3. Lack of Support for Data Preparation and External Constraints
Participants in our study shared various challenges with data gathering, consolidation, featurization (e.g., curating new hypothesized drivers for KPI goals), aggregation from multiple sources, and cleaning it to work around missing, inaccurate, and outlier data. These operations are already hard for expert analysts (Dimara et al., 2021; Crisan and Fiore-Gartland, 2021; Kandel et al., 2012; Alspaugh et al., 2019), and our study confirms that it can be all the more daunting for business users. For example, the participants raised some fundamental issues with adding qualitative data (procured over calls, meetings, chats, and reviews) as drivers into models, treating connected drivers (e.g., calls and meetings), and resolving ambiguities about the drivers or KPIs as they could mean different things to different teams at times. Therefore, participants suggested co-designing and tight integration of WIA techniques with data wrangling and preparation systems.
@@
Further, participants also wanted to factor in economical (e.g., inflation, fed rate hike) and social (e.g., COVID, war, street protests) constraints into advanced analytics (P6, P9, P10, P15) since they had very significant impacts on businesses. Participants often dealt with these constraints for making decisions by turning to their intuition (Wang et al., 2019) and domain expertise. As a consequence, participants indicated that WIA techniques need to be enhanced to increase their reliability so that it can actively be used “as an adjunct” (P8) to inform their decisions. Therefore, future tools need to focus on feedback mechanisms between business users and advanced analytics which will allow users’ expertise as well as external constraints to be easily integrated into predictions (Zhang et al., 2020; Honeycutt et al., 2020).
@@ 8. Discussion
Our two-part study confirms that business users are eager to independently conduct data-driven advanced analytics to inform their business decisions. Based on our findings, we discuss design recommendations for future business analytics systems to better support these advanced analytics needs.
@@ 8.1. Enhance Interpretation and Confidence
Our study reveals that business users are far from naive; they possess strong domain expertise and common sense, enabling them to critically assess predictions from automated tools. All participants indicated that WIA predictions cannot be followed blindly (Wang et al., 2019; Kaur et al., 2024), recognizing that even the most advanced models are based on limited data. This data often fails to capture all relevant constraints and domain-specific knowledge, resulting in predictions that may not be entirely accurate. However, participants emphasized the value of advanced analytics in complementing their decision-making, supporting their actions with data-driven evidence.
To foster trust and ensure users feel confident using data-driven analyses, systems must not only be transparent in communicating confidence levels in predictions but also offer actionable steps to calibrate these confidence levels. Participants suggested that model accuracy scores or error metrics should be corroborated with risk factors that communicate potential negative outcomes, such as potential financial losses or the emergence of new competitors or products. Moreover, users desired guidance on next steps based on risk levels (e.g., a medium risk factor might suggest gathering more data and re-running the analysis). Recent work has explored general users’ perceptions of trust and performance metrics (Langer et al., 2022; Zhang et al., 2020; Honeycutt et al., 2020; Bhattacharya et al., 2024) Conducting similar studies specifically with business users can reveal other factors that enhance their trust in model predictions and, in turn, guide the design of advanced analytics systems that better meet their needs. Additionally, future research should investigate the requirements of technically proficient business users, focusing on analyses features that optimize their decision-making. This opens up a substantial opportunity for interdisciplinary collaboration to refine WIA techniques, making them more relevant and effective for business contexts.
@@ 8.2. Foster Communication and Team Collaboration
Participants consistently emphasized the importance of WIA in empowering business users to justify their decisions with data, providing them with a stronger voice when navigating organizational hierarchies. As business decisions often require approval from executive stakeholders, it is crucial for users to effectively communicate the rationale behind the proposed decision action predictions of advanced analytics techniques, especially since neither the business users nor the executives are typically data science or machine learning experts. Since business decisions often require approval from executive stakeholders, it is essential for users to clearly communicate the rationale behind decision-making predictions generated by advanced analytics techniques. Additionally, this must be done with the understanding that both business users and executives usually lack expertise in data science or machine learning. This presents a clear opportunity to rethink traditional tools used. For example, current dashboards typically display static summaries, such as the percentage of goals achieved or fixed drill-down paths into data segments (Tory et al., 2021; Sarikaya et al., 2018). These high-level overviews, while useful, do not provide the flexibility needed for decision-making contexts (Oral et al., 2023). Similarly, spreadsheet tables can be cumbersome and confusing (Bartram et al., 2021), and individual presentations prepared for every meeting require substantial manual effort. By contrast, interactive visual tools accommodating advanced analytics could revolutionize these processes. Instead of static, one-way communication, such tools could allow users to dynamically explore various scenarios and their impacts on KPIs. For example, during a meeting, a business user could adjust variables in real-time to demonstrate how different marketing strategies might affect sales forecasts. These tools could facilitate live discussions and decision-making, similar to how participants could see the immediate outcomes of various strategies. Moreover, these tools could be designed to quickly incorporate feedback from executives or other team members during the meeting itself (Spitzeck and Hansen, 2010), rather than requiring additional time to go back, rerun analyses, and return with new findings. For instance, if an executive recalls a similar scenario from past experience or suggests adjusting a particular variable due to new constraints, the tool could instantly update the analysis to reflect that feedback. This kind of interactive, iterative process would not only save time but also foster collaboration, ensuring that decisions are informed by both data and experiential knowledge without the delays that are common in current workflows. Further, emerging technologies such as natural language interfaces and interactive storytelling platforms (Sun et al., 2024; He et al., 2024) must also be leveraged in such tools to go beyond merely presenting predictive results to also additionally explain, modify, and contextualize the predictions from advanced analytics.
@@ 8.3. Scale WIA with Decision Management Support
While WIA techniques allow for rapid computation of what-if scenarios and offer insights into various decision paths, participants observed that the number of scenarios and corresponding decision paths can grow exponentially as the number of driver variables and constraints increase. Hence, managing and reusing these analyses at scale presents a formidable challenge, especially as users increasingly gain access to real-time data. To prevent business users from becoming overwhelmed or, as one participant described, from “getting lost in the sauce” (P9), it is essential to develop robust decision management systems that can effectively monitor, track, and log users’ analyses. These systems must be capable of handling large volumes of data and complex decision paths without compromising usability. Though such provenance-based systems have been widely developed for software developers (Schreiber et al., 2021; Packer et al., 2019), data analysts (Cutler et al., 2020; Madanagopal et al., 2019; Ragan et al., 2015), data scientists (Namaki et al., 2020; Souza et al., 2022; Kumar et al., 2016), and researchers (Jun et al., 2019, 2022a, 2022b; Gu et al., 2023) for various analyses (e.g., visual analysis, statistical analysis, software development, etc.), there is gap in supporting non-technical domain users like business professionals. These users require more accessible and user-friendly solutions tailored specifically for analyses that help inform data-driven decisions.
@@
Further, decision management support systems can significantly enhance incorporation of WIA by enabling tools for ranking, comparing, and recommending optimized decision paths. Our study indicates that such integration, as suggested by business users, will significantly improve their ability to leverage advanced analytics for truly data-driven decision-making. However, to maximize the success of these systems, it is essential to focus on usability and accessibility. For instance, systems supporting multitude of models, tons of features, and complex visualizations and workflows or programming language-based tools will not meet the needs of business users. While natural-language, query-based features powered by LLMs may be very useful, as seen in software testing (Santos et al., 2024) and cyber-physical systems (Elmaaroufi et al., 2024). Therefore, besides the vast scope for machine learning modeling and optimization research, there is a huge opportunity for user experience and data and visual analytics communities to contribute to the development of scalable, user-friendly decision management systems that effectively support advanced analyses for decision-making.

### v4

@@ 7.2.3. Streamlines Decision-Making and Enhances Teamwork
Almost all participants agreed that their existing analysis processes could not be replicated easily to justify their decision actions. For instance, multiple participants commented that their current decision actions were based on “trial and error” (P3, P14, P19, P21), which they often deemed to be inefficient and unpredictable, making it challenging to rely on. Recognizing the limitations of their current practices, participants saw the potential of WIA to streamline decision processes, making them more accessible, reproducible, and actionable. This consequently allowed for “consistent” (P3, P10) (Kandel et al., 2012) and “smoother” (P22) decision-making. Further, in addition to WIA techniques helping them in their own analysis, participants also added that it will help facilitate collaborative analyses for decision-making envisioning a scenario where stakeholders could engage in decision reasoning collectively. For example, one participant P4 mentioned how they could “take [this] to their job today without the support of IT… also give it to other team members”. Therefore, participants explained that advanced analytics techniques would help them make robust decisions as they could get external perspectives and benefits from others’ experiences and expertise through standardized and collaborative analysis processes.
@@ 7.3. Opportunities in Enhancing the Use of WIA for Making Data-Driven Decisions
7.3.1. Currently Inaccessible to Business Professionals
Many participants like P1, P7, P8 mentioned that even though they had heard of features similar to some representative WIA techniques in existing commercial BI tools (e.g., key influencers in Power BI, importance and correlation analysis in Einstein, etc.), they had “no idea or not know too much [about them] or barely used them”. Upon further discussion, participants revealed that the high costs associated with such tools rendered them inaccessible. Additionally, participants voiced concerns about the complexity of these tools, which typically offered a plethora of models and parameters to navigate, or deemed them to be too general and inefficient to use for their own use cases. The overwhelming number of features too made it difficult for professionals to effectively use the tools, a challenge exacerbated by automatically generated results that lacked clear explanations for them to draw meaningful conclusions. Moreover, participants cautioned the requirement of technical expertise to utilize these tools effectively, a skill set that many business professionals did not possess. Furthermore, business professionals also lacked the time and incentives to learn or to keep up with such technologies, which are rapidly evolving.
Despite their enthusiasm for adopting WIA in their work, it seems that tools are not currently designed to meet their needs. The vast space of possible scenarios seems to be a major problem in particular. P15, for instance, frustratingly mentioned how they have “…to [currently] commit to just solving the first thing… that is going to be the most bang for the buck…. [because they]… have 900 problems happening all at once…”. Such expressions inform us that business professionals have no time to experiment with all possible scenarios but need flexibility to quickly identify the first set of actionable decisions that yield favorable business outcomes and play with them to account for constraints, a concept akin to “satisficing” (Simon, 1960), suggesting that more complex analyses atop WIA, like recommendations or ranking of WIA scenarios, could help.
@@ 7.3.2. Difficulties in Interpreting and Understanding Predictions
While participants expressed enthusiasm about using WIA techniques for decision-making, they emphasized that analytics alone cannot make the decisions. Their domain expertise remains critical for developing effective strategies (Wang et al., 2019). For instance, P4 pointed to how “common human understanding may not translate to the model without [their] external involvement”, giving an example where an increase in unit price beyond a threshold will decrease demand, thereby decreasing sales rather than increasing it, a nuance that may not be fully captured by models. Participants emphasized the need for transparency in how predictions were generated and wanted clarification on when and to what extent to adopt predictions. For instance, they wanted to know how to handle “unrealistic [predictions]” (P5, P6, P22), or when “so many factors [or decision paths]” (P2) are shown to them, and how to “[include] factors that cannot be captured in the data” (P2, P18).
Another recurring challenge was interpreting the accuracy of predictions (low v/s fair v/s high) and translating this into confidence in next steps (Elhamdadi et al., 2022). For example, should low-scored models be treated as worst-case scenarios and perform additional manual analysis to make decisions? Or should they take medium scored models as baseline and expect risks in lower scored models? While senior, more experienced business professionals understood that predictions should be taken with caution, they emphasized the importance of training junior professionals not to follow predictions blindly. Addressing these concerns requires WIA tools to foster deeper transparency and provide interpretable predictions that can help enhance business professionals’ trust in the predictions.
@@ 7.3.3. Lack of Support for Data Preparation and External Constraints
Participants reported various challenges in data gathering, consolidation, featurization (e.g., curating new hypothesized drivers for KPI goals), aggregation data from multiple sources, and cleaning it to work around missing, inaccurate, and outlier data. These tasks are already complex for expert analysts (Dimara et al., 2021; Crisan and Fiore-Gartland, 2021; Kandel et al., 2012; Alspaugh et al., 2019), and our study confirms that it can be even more daunting for business professionals. For example, the participants highlighted issues such as incorporating qualitative data (e.g., meetings, chats, and reviews) as drivers into models, handling interdependent drivers (e.g., calls and meetings), and resolving ambiguities related to the drivers or KPIs, which often carry different meanings across teams. To address these issues, participants suggested co-designing WIA techniques with seamless integration into data wrangling and preparation systems.
@@
Further, participants also wanted to factor in economical (e.g., inflation, fed rate hike) and social (e.g., COVID, war, street protests) constraints into advanced analytics (P6, P9, P10, P15) since they had very significant impacts on businesses. Participants often dealt with these constraints for making decisions by turning to their intuition (Wang et al., 2019) and domain expertise. As a consequence, participants indicated that WIA techniques need to be enhanced to increase their reliability so that it can actively be used “as an adjunct” (P8) to inform their decisions. Therefore, future tools need to focus on feedback mechanisms between business professionals and advanced analytics which will allow professionals’ expertise as well as external constraints to be easily integrated into predictions (Zhang et al., 2020; Honeycutt et al., 2020).
@@ 8. Discussion
Our two-part study confirms that business professionals are eager to independently conduct data-driven advanced analytics to inform their business decisions. Based on our findings, we discuss design recommendations for future business analytics systems to better support these advanced analytics needs.
@@ 8.1. Enhance Interpretation and Confidence
Our study reveals that business professionals are far from naive; they possess strong domain expertise and common sense, enabling them to critically assess predictions from automated tools. All participants indicated that WIA predictions cannot be followed blindly (Wang et al., 2019; Kaur et al., 2024), recognizing that even the most advanced models are based on limited data. Data often fails to capture all relevant constraints and domain-specific knowledge, resulting in predictions that may not be entirely accurate. Even commercial tools like Power BI (BI, 2018), which provide advanced analytics features such as correlation insights or key influencers (akin to driver importance analysis), present these outputs to professionals but leave the responsibility of assessing their trustworthiness to the users. Participants emphasized the value of advanced analytics in complementing their decision-making and supporting their actions with data-driven evidence. However, they emphasized the need for systems that foster trust and confidence in the predictions they produce.
To achieve this, systems must go beyond merely displaying confidence levels buried within multiple features of commercial tools. They should transparently communicate prediction reliability and offer actionable steps to help professionals calibrate and validate these confidence levels effectively. Participants suggested that model accuracy scores or error metrics should be corroborated with risk factors that communicate potential negative outcomes, such as potential financial losses or the emergence of new competitors or products. Moreover, professionals desired guidance on next steps based on risk levels (e.g., a medium risk factor might suggest gathering more data and re-running the analysis). Several user studies (Langer et al., 2022; Zhang et al., 2020; Honeycutt et al., 2020; Bhattacharya et al., 2024) have explored users’ perceptions of trust and performance metrics in tasks such as model steering and predicting income group. Conducting similar studies specifically with business professionals for what-if analysis tasks can reveal other factors that enhance their trust in model predictions and, in turn, guide the design of advanced analytics systems that better meet their needs. Additionally, future research should investigate the requirements of technically proficient business professionals, focusing on analyses features that optimize their decision-making. This opens up a substantial opportunity for interdisciplinary collaboration to refine WIA techniques, making them more relevant and effective for business contexts.
@@ 8.2. Foster Communication and Team Collaboration
Participants consistently emphasized the importance of WIA in empowering business professionals to justify their decisions with data, providing them with a stronger voice when navigating organizational hierarchies. Since business decisions often require approval from executive stakeholders, it is essential for professionals to clearly communicate the rationale behind decision-making predictions generated by advanced analytics techniques. Additionally, this must be done with the understanding that both business professionals and executives usually lack expertise in data science or machine learning. This presents a clear opportunity to rethink traditional tools used.
@@
For example, current dashboards like Tableau (Tableau, 2024) or Power BI (BI, 2018) typically display static textual summaries, such as the percentage of goals achieved or fixed drill-down paths into data segments (Tory et al., 2023; Sarikaya et al., 2018). These high-level overviews, while useful, do not provide the flexibility needed for decision-making contexts (Oral et al., 2023). Similarly, tables created in spreadsheet-based tools can be cumbersome and confusing (Bartram et al., 2021), and cannot be modified on the fly during discussions or meetings. Further, translating insights from tables to individual presentations for every meeting requires substantial manual effort. By contrast, interactive visual tools accommodating advanced analytics could revolutionize these processes. Instead of static, one-way communication, such tools could allow professionals to dynamically explore various scenarios and their impacts on KPIs. For example, during a meeting, a business professional could adjust variables in real time to demonstrate how different marketing strategies may affect sales forecasts. These tools could facilitate live discussions and decision-making, as participants could see immediate outcomes of various strategies. In addition, these tools could be designed to quickly incorporate feedback from executives or other team members during the meeting itself (Spitzeck and Hansen, 2010), rather than requiring additional time to go back, rerun analyses and return with new findings. For instance, if an executive recalls a similar scenario from past experience or suggests adjusting a particular variable due to new constraints, the tool could instantly update the analysis to reflect that feedback. This kind of interactive, iterative process would not only save time but also foster collaboration, ensuring that decisions are informed by both data and experiential knowledge without the delays that are common in current workflows.
Further, emerging technologies such as natural language interfaces, interactive storytelling platforms (Sun et al., 2024; He et al., 2024), and AI assistants (Perez, 2024; et al., 2024) must also be leveraged in such tools to go beyond merely presenting predictive results to also additionally explain, modify, and contextualize the predictions from advanced analytics.
@@ 8.3. Scale WIA with Decision Management Support
While WIA techniques allow for rapid computation of what-if scenarios and offer insights into various decision paths, participants observed that the number of scenarios and corresponding decision paths could grow exponentially (Liu et al., 2020a, b) as the number of driver variables and constraints increase. Hence, managing and reusing these analyses at scale presents a formidable challenge, especially as professionals increasingly gain access to real-time data. To prevent business professionals from becoming overwhelmed or, as one participant described, from “getting lost in the sauce” (P9), it is essential to develop robust decision management systems that can effectively monitor, track, and log their analyses.
Additionally, such systems can also support professionals in formalizing their decision-making processes. For instance, they can assist in clearly identifying goals, drivers, and constraints, which is a critical step in structuring decisions and aligning them with business objectives. By providing this structure, future systems can be developed with targeted technological support to address the complex decision paths and associated challenges (Khadpe et al., 2024; Dimara et al., 2021). For instance, these systems must be capable of handling large volumes of data and complex decision paths without compromising usability. Though such provenance-based systems have been widely developed for software developers (Schreiber et al., 2021; Packer et al., 2019), data analysts (Cutler et al., 2020; Madanagopal et al., 2019; Ragan et al., 2015), data scientists (Namaki et al., 2020; Souza et al., 2022; Kumar et al., 2016), and researchers (Jun et al., 2019, 2022a, 2022b; Gu et al., 2023) for various analyses (e.g., visual analysis, statistical analysis, software development, etc.), there is gap in supporting domain experts like business professionals with accessible tools for using advanced analytics (Tory et al., 2023; Dimara et al., 2021). These professionals require more user-friendly solutions tailored specifically for analyses that help inform data-driven decisions. For instance, professionals should be able to specify their hypotheses and parameters effortlessly without any need of technical expertise (Bartram et al., 2021). Hence, we develop our probe as an intuitive interactive interface requiring minimal external guidance. The tools should also support the seamless incorporation of domain knowledge and context, allowing professionals to leverage their expertise directly within the analysis, similar to how we allow the constraints to be specified. Furthermore, fostering trust in the models is essential, so tools must be designed to enhance transparency and make interpretations more accessible (Wang et al., 2019; Yang et al., 2020; Elhamdadi et al., 2022). Finally, these solutions should simplify the management and comparison of multiple scenarios, enabling professionals to evaluate alternatives and make actionable decisions with greater ease and efficiency. By integrating formalization into decision management systems, these tools can bridge the gap between sophisticated analytics and the practical demands of business decision-making.
@@
Further, decision management support systems can significantly enhance incorporation of WIA by enabling tools for ranking (Gratzl et al., 2013; Pajer et al., 2017), comparing, and recommending optimized decision paths (Liu et al., 2020b). Our study, along with others  (Dimara et al., 2021; Dimara and Stasko, 2021; Tory et al., 2023) indicate that such integration, especially as suggested by business professionals, will significantly improve their ability to leverage advanced analytics for truly data-driven decision-making. However, to maximize the success of these systems, it is essential to focus on usability and accessibility. For instance, neither programming-based tools nor systems supporting a multitude of models, features, complex visualizations, and workflows will meet the needs of business professionals. Beyond supporting WIA for structured tabular data, business use cases frequently involve qualitative data, such as expert opinions, contextual knowledge, or industry trends. Incorporating these unstructured data types into the decision-making process is also essential, as they provide critical context and insights that complement quantitative analyses. In this context, natural-language, query-based features powered by LLMs offer significant potential. These features have already demonstrated their utility in domains such as software testing (Santos et al., 2024) and cyber-physical systems (Elmaaroufi et al., 2024), and could similarly enable business professionals to interact with complex advanced analytics and incorporation of textual data within it more intuitively. Therefore, besides the vast scope for machine learning modeling and optimization research, there is a huge opportunity for user experience and data and visual analytics communities to contribute to the development of scalable, user-friendly decision management systems that effectively support advanced analyses for decision-making.